OpenStudy (anonymous):

divide and simplify 4x/7 divided by 16x+32/7x+17

OpenStudy (anonymous):

invert the second function and multiply 4x/7 X 7x+17/ 16x+32 = 4x(7x+17) / 7(16x+32) = 28x^2 +68x / 112x + 224 = 4x(7x + 17) / 112(x+2) = x(7x + 17 )/ 28(x+2)